OpenStudy (anonymous):

College Algebra Question:

OpenStudy (anonymous):

So I know how to do both, its just giving an answer about how they are similar is what has me stumped

OpenStudy (freckles):

hint: conjugate

OpenStudy (anonymous):

That hint doesn't really help

OpenStudy (freckles):

\[\text{ simplify } \frac{1}{\sqrt{2}-3} \text{ and also simplfy } \frac{1-i}{2+i}\] not sure how it doesn't since that is what you to need to simplify both of these you need to multiply both top and bottom by the conjugate of the bottom

OpenStudy (anonymous):

okay im working it out give me a sec

Nnesha (nnesha):

\[\large\rm \frac{ 1-i }{ 2+i } \] example #1 \[\frac{ 1 }{ \sqrt{2}-3 }\] ex #2 nvm solve freckless example :-)

OpenStudy (anonymous):

okay @ freckles first problem is \[\sqrt{2}+3\]

OpenStudy (anonymous):

@freckles and the second is \[\frac{ 1 }{ 5 }-\frac{ 3 }{ 5 }i\]

Nnesha (nnesha):

what should be at the denominator ?

Nnesha (nnesha):

first one is not right \[\frac{ 1 }{ \sqrt{2}-3 }\] use freckles hint: multiply by conjugate

Nnesha (nnesha):

top part is right you should multiply both numerator and denominator by conjugate

OpenStudy (anonymous):

now I got \[-\frac{ 1 }{ 7 }\]

Nnesha (nnesha):

\[\huge\rm \frac{ 1 }{ \sqrt{2}-3} \times \frac{\sqrt{2}+3}{\sqrt{2}+3}\] multiply denominator by denominator and numerator by numerator so 1 times {sqrt{2}+3} = sqrt{2}+3 that should stay at the top and yes you are right denominator is -7

Nnesha (nnesha):

so what was ur first step for both questions ??

OpenStudy (anonymous):

So the first step would be to conjugate the denominator and then FOIL.. ( am i using the right terms lol)

Nnesha (nnesha):

right first step multiply by conjugate

OpenStudy (anonymous):

Okay great I get it now @Nnesha

Nnesha (nnesha):

great

Nnesha (nnesha):

freckles is online you can ask question :-)

OpenStudy (freckles):

yeah my internet has been wacky lately

OpenStudy (anonymous):

Hey @freckles so as i write my paragraph to explain how they are similar im first going to state how you have to conjugate the denominator. Then you multiply across for both numerator and denominator would that be considered the foil method or am i using the wrong term?

OpenStudy (freckles):

I would just call it multiplying stead of foil. For something of the forms: \[\frac{c+d}{\sqrt{s}+\sqrt{t}} \text{ and } \frac{c+di}{s+ti}\] to simplify you need to first identify the conjugate of the denominator once you have identify the conjugate of the denominators, you will take that conjugate and just multiply your fraction on both top and bottom by that conjugate.

OpenStudy (freckles):

I don't like foil method :p though most people say to multiply something in the form (a+b)(c+d) you can do that by foil so it wouldn't be totally wrong to say it, it just isn't very mathematical to say but I'm sure your teacher would be pleased either way since you know I think the point of the exercise is for you to understand and for you to come up with your own explanation so I think your explanation would be acceptable

OpenStudy (anonymous):

lol okay

Nnesha (nnesha):

i guess you can say foil it bec \[\frac{ c+d }{ \sqrt{s}+\sqrt{d} } \times \frac{\sqrt{s}-\sqrt{t} }{ \sqrt{s}-\sqrt{t} }\]  if you write it as one fraction u r using foil method